The formula h= 16t2 represents the distance h; infeet, that an object falls from rest after 1 seconds. In how many seconds will an object fall 400 ft.

Answer by checkley77(12844)

400=16t^2
t^2=400/16
t^2=25
t=sqrt25
t=5 seconds.